# Electric field induction in quark-gluon plasma due to thermoelectric effects

### v1

@@ 4.2 Induced electric field in anisotropic system
5 Summary and Conclusion
In summary, we have calculated thermoelectric coefficients for evolving QGP medium created in RHICs. Thermoelectric transport phenomena can induce an electric field in the medium. We have estimated the same for the first time using Gubser hydrodynamics. We used the kinetic theory-based RTA approach to calculate thermoelectric coefficients. Unlike the static picture of QGP considered in all the earlier calculations Das2021 ; Abhishek:2020wjm ; Kurian:2021zyb ; PhysRevD.102.096011 , we solved the Boltzmann equation for evolving QGP by taking care of temperature evolution. Deriving the general expressions of the transport coefficients, we used different cooling rates (ideal hydrodynamic, ideal MHD, Bjorken flow) for their estimation. We used a quasiparticle model to reproduce the lattice QCD EoS for QGP. Moreover, calculations and results are obtained for head-on and peripheral collisions where we expect a magnetic field due to spectator currents. All the results are at finite baryon chemical potential $\mu_{B}=0.3$ GeV.
@@ 5 Summary and Conclusion
In a fluid of charged particles with finite thermal conductivity, thermoelectric phenomena can create or induce an electromagnetic field. QGP consists of quarks, which are electrically charged and have finite baryon quantum numbers. Therefore, they are capable of electrical (charge) and thermal (baryon) conduction. As a result, thermoelectric transport in QGP at finite baryon chemical potential leads to the induction of an electric field. QGP medium is governed by the QCD interaction. Studies show that the EM field of the order of $\Lambda_{\rm QCD}$ scale can interfere with the interactions, contributing to many interesting phenomena PhysRevD.78.074033 ; PhysRevLett.73.3499 ; PhysRevD.56.5251 . It is widely accepted that such a huge EM field was created in peripheral collisions Tuchin2013ie . This work showed that an EM field can also be created during the QGP evolution due to the thermoelectric effect, even in head-on collisions. The maximum electric field produced in head-on and peripheral collisions is $\sim 0.6~{}m_{\pi}^{2}$ and $0.9~{}m_{\pi}^{2}$, respectively, which are two to one times smaller than the QCD scale ($\sim 2~{}m_{\pi}^{2}$). However, this can contribute to many interesting phenomena, as mentioned before. Nevertheless, the impact might be small. It is essential to mention here that we have taken care of temperature evolution in calculating thermoelectric coefficients. i.e., the effect of cooling is considered, which was missing in all the earlier studies. Results for static and evolving pictures are compared, and we found that the impact of evolution is significant though non-monotonic. In peripheral collisions, we considered an external time-varying magnetic field. The effect of the initial magnetic field and its decay parameter are also studied. The induced electric field falls off with time. At the early time of evolution, the magnetic decay parameter has a significant impact; however, the strength of the initial magnetic field has a minor effect. Considering RHIC and LHC matter, we considered the external magnetic field along the $y$ direction for the numerical estimation. However, we have also generalized the calculations, where the direction of the magnetic field was not preferred, and we obtained the thermoelectric coefficient matrix for the same.
@@ A.1 Case-I (B = 0)
To find $\delta f_{i}$ in Eq. (15), we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA). In the presence of temperature gradients, the BTE under RTA can be expressed as PhysRevD.102.016016
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. $\vec{E}$ is an external electric field. We can solve Eq. (A.1) using Eqs. (18) and (19), and get the expression of $\alpha_{j}$’s. Here, the first term in left-hand side (lhs) of the Eq. (A.1) becomes,
$-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}{\dot{\vec{E}}}+\dot{\alpha_{2}}{\vec{\nabla}T}+\alpha_{2}{\vec{\nabla}\dot{T}}\Big{\}}.$ (A.2)
The second term in lhs of the Boltzmann equation leads to,
$\frac{\partial f_{i}}{\partial x_{i}}$ $=\frac{\partial}{\partial x_{i}}(f^{0}_{i}+\delta f_{i})=\frac{\partial}{\partial x_{i}}(f^{0}_{i})+\frac{\partial f_{i}}{\partial x_{i}}(\delta f_{i})=\frac{\partial}{\partial x_{i}}\Big{(}\frac{1}{1\pm\exp{\big{(}\beta(\omega_{i}-{\rm b}_{i}\mu_{i})\big{)}}}\Big{)}+0$
@@
The third term in lhs leads to,
@@
$-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\omega_{i}{\vec{v}_{i}}.\Big{(}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}{\dot{\vec{E}}}+\dot{\alpha_{2}}{\vec{\nabla}T}+\alpha_{2}{\vec{\nabla}\dot{T}}\Big{)}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}-(\omega_{i}-{\rm b}_{i}h)\frac{\vec{\nabla}T}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}$
$+q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}=-\frac{\omega_{i}}{\tau_{R}^{i}}\Big{(}\alpha_{1}\vec{E}+\alpha_{2}\vec{\nabla}T+\alpha_{3}\vec{\nabla}\dot{T}\Big{)}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.6)
@@
$\dot{\alpha_{1}}$ $=-\Bigg{(}\frac{1}{\tau_{R}^{i}}\alpha_{1}+\frac{q_{i}}{\omega_{i}}\Bigg{)},~{}~{}~{}\dot{\alpha_{2}}=-\bigg{(}\frac{1}{\tau_{R}^{i}}\alpha_{2}-\frac{\omega_{i}-{\rm b}_{i}h}{\omega_{i}T}\bigg{)},~{}~{}~{}{\alpha_{3}}=-{\tau_{R}^{i}}\alpha_{2},$ (A.7)
@@
From Eq. (A.12) and Eq. (A.8) we can see that $Y_{p}$ is a column matrix with the same coefficients as that of matrix X. Further, the differentiation of Eq. (A.12) with respect to time gives us $Y_{p}^{{}^{\prime}}$ = $Y^{{}^{\prime}}U+YU^{{}^{\prime}}$, where $Y^{{}^{\prime}}=AY$. Hence, $Y_{p}^{{}^{\prime}}$ = $AY_{p}+YU^{{}^{\prime}}$. Comparison of above equation with Eq. (A.8) tells us
@@
$\alpha_{1}=-\frac{q_{i}\tau_{R}^{i}}{\omega_{i}},~{}~{}~{}\alpha_{2}=\frac{\tau_{R}^{i}(\omega_{i}-{\rm b}_{i}h)}{T\omega_{i}},~{}~{}~{}\alpha_{3}=-\tau_{R}^{i}\alpha_{3}.$
@@
$\delta f_{i}$ $=-q_{i}\tau_{R}^{i}(\vec{v_{i}}.\vec{E})+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T}\Big{[}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}-\tau_{R}^{i}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.18)
Here $\pm$ indicates a positive sign for positively charged particles(antiparticles) and a negative sign for negatively charged particles (antiparticles).
@@ A.2 Case-II (B $\not$= 0)
To find the expression of $\delta f_{i}$ in Eq. (15), we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA). In the presence of an external electromagnetic field, BTE under RTA can be expressed as PhysRevD.102.016016
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. We can solve Eq. (A.19) using Eqs. (38) and (4.1.1), and get the expression of $\alpha_{j}$’s. Here, the first term in left-hand side (lhs) of the Eq. (A.19) becomes,
$\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}\dot{{\vec{E}}}+\dot{\alpha_{3}}{\vec{\nabla}T}+{\alpha_{3}}{\vec{\nabla}\dot{T}}+\alpha_{5}({\vec{\nabla}T}\times\dot{{\vec{B}}})+\dot{\alpha_{5}}({\vec{\nabla}T}\times{\vec{B}})+{\alpha_{5}}({\vec{\nabla}\dot{T}}\times{\vec{B}})$
$+\alpha_{5}(\vec{\nabla}\dot{T}\times{\vec{B}})+\alpha_{8}({\vec{E}}\times\dot{{\vec{B}}})+\dot{\alpha_{8}}({\vec{E}}\times{\vec{B}})+\alpha_{8}(\dot{\vec{E}}\times{\vec{B}})-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.20)
The second term in lhs of the Boltzmann equation leads to,
$\frac{\partial f_{i}}{\partial x_{i}}$ $=\frac{\partial}{\partial x_{i}}(f^{0}_{i}+\delta f_{i})=\frac{\partial}{\partial x_{i}}(f^{0}_{i})+\frac{\partial f_{i}}{\partial x_{i}}(\delta f_{i})=\frac{\partial}{\partial x_{i}}\Big{\{}\frac{1}{1\pm\exp{\big{(}\beta(\omega_{i}-{\rm b}_{i}\mu_{i})\big{)}}}\Big{\}}+0$
@@
The third term in lhs leads to,
@@
$q_{i}\left(\vec{E}+\frac{\vec{k}_{i}}{\omega_{i}}\times\vec{B}\right)\cdot\frac{\partial f_{i}}{\partial\vec{k_{i}}}=q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\big{(}\vec{v}\times\vec{B}\big{)}\big{\{}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})$
$+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}})\big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.23)
@@
$\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}\dot{{\vec{E}}}+\dot{\alpha_{3}}{\vec{\nabla}T}+{\alpha_{3}}{\vec{\nabla}\dot{T}}+\alpha_{5}({\vec{\nabla}T}\times\dot{{\vec{B}}})+\dot{\alpha_{5}}({\vec{\nabla}T}\times{\vec{B}})+{\alpha_{5}}({\vec{\nabla}\dot{T}}\times{\vec{B}})$
$+\alpha_{5}(\vec{\nabla}\dot{T}\times{\vec{B}})+\alpha_{8}({\vec{E}}\times\dot{{\vec{B}}})+\dot{\alpha_{8}}({\vec{E}}\times{\vec{B}})+\alpha_{8}(\dot{\vec{E}}\times{\vec{B}})-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$
$-(\omega_{i}-{\rm b}_{i}h)\vec{v_{i}}\cdot\frac{\vec{\nabla}T}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\big{\{}-\alpha_{1}q_{i}\vec{v_{i}}\cdot(\vec{E}\times\vec{B})-\alpha_{2}q_{i}\vec{v_{i}}\cdot(\dot{\vec{E}}\times\vec{B})$
$-\alpha_{3}q_{i}\vec{v_{i}}\cdot(\vec{\nabla}T\times\vec{B})-\alpha_{4}q_{i}\vec{v_{i}}\cdot(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{5}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\vec{\nabla T}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{\nabla}T)({\vec{B}}\cdot\vec{B})\}$
@@
$+\alpha_{8}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\vec{E}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{E})({\vec{B}}\cdot\vec{B})\}+\alpha_{9}q_{i}\{-(\vec{v_{i}}\cdot\dot{\vec{B}})(\vec{E}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{\dot{E}})({\vec{B}}\cdot\dot{\vec{B}})\}$
$+\alpha_{9}q_{i}\{-(\vec{v_{i}}\cdot\dot{\vec{B}})(\vec{E}\cdot\vec{B})+(\vec{v_{i}}\cdot\dot{\vec{E}})({\vec{B}}\cdot\dot{\vec{B}})\}+\alpha_{10}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\dot{\vec{E}}\cdot\vec{B})+(\vec{v_{i}}\cdot\dot{\vec{E}})({\vec{B}}\cdot\vec{B})\}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}},$ (A.24)
@@
$=-\frac{\omega_{i}}{\tau_{R}^{i}}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})$
$+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.25)
@@
Here, $\dot{\alpha_{1}}$, $\dot{\alpha_{3}}$, $\dot{\alpha_{5}}$ and $\dot{\alpha_{5}}$ from Eq. (A.2) can be expressed in terms of matrix equation as
@@
$X=\begin{pmatrix}\alpha_{1}\\ \alpha_{3}\\ \alpha_{5}\\ \alpha_{8}\end{pmatrix},~{}~{}~{}A=\begin{pmatrix}-\frac{1}{\tau_{R}^{i}}&0&0&-\frac{q_{i}F^{2}}{\omega_{i}}\\ 0&-\frac{1}{\tau_{R}^{i}}&-\frac{q_{i}F^{2}}{\omega_{i}}&0\\ 0&\frac{q_{i}}{\omega_{i}}&-\frac{1}{\tau_{R}^{i}}&0\\ \frac{q_{i}}{\omega_{i}}&0&0&-\frac{1}{\tau_{R}^{i}},\end{pmatrix},~{}~{}~{}G=\begin{pmatrix}-\frac{q_{i}}{\omega_{i}}\\ \frac{\omega_{i}-{\rm b}_{i}h}{T\omega_{i}}\\ 0\\ 0\end{pmatrix},$

### v2

@@ IV.2.1 Significance of the findings
V Summary and Conclusion
In summary, for the first time, we estimated the induced electric field in QGP solely using the thermoelectric effect. We employed the kinetic theory-based RTA approach to calculate thermoelectric coefficients and induced field. For numerical estimation, we used a quasiparticle model that reproduces the lattice QCD EoS of QGP. An analytical expression of the cooling rate is necessary to calculate the induced electric field, for which we use Gubser hydrodynamics. Moreover, considering the realistic scenario of the temperature evolution of QGP, we introduced a term containing the time derivative of temperature gradient in the calculation of the total single-particle distribution function in the Boltzmann equation. We named it the evolving picture, and the scenario where it is not considered is termed a static picture. We have also studied their relative significance. Heat conduction demands a conserved charge (or particle number). Here, we have baryon number conservation. In this instance, a finite baryon chemical potential is necessary. Otherwise, thermoelectric coefficients diverge up at vanishing baryon density. All results here are presented for the baryon chemical potential $\mu_{B}=0.3$ GeV.
@@ V Summary and Conclusion
In a fluid of charged particles with finite thermal conductivity, thermoelectric phenomena can create or induce an electromagnetic field. QGP consists of quarks, which are electrically charged and have finite baryon quantum numbers. Therefore, they are capable of electrical (charge) and thermal (baryon) conduction. As a result, thermoelectric transport in QGP at finite baryon chemical potential leads to the induction of an electric field. QGP medium is governed by the QCD interaction. Studies show that the EM field of the order of $\Lambda_{\rm QCD}$ scale can interfere with the interactions, contributing to many interesting phenomena [72, 28, 73]. It is widely accepted that such a huge EM field was created in peripheral collisions [53]. The present work shows that an EM field can also be created during the QGP evolution due to the thermoelectric effect, even in head-on collisions. The maximum electric field produced in head-on and peripheral collisions is $\sim 0.7~{}m_{\pi}^{2}$ and $1.0~{}m_{\pi}^{2}$, respectively, which are close to the QCD scale ($\sim 1m_{\pi}^{2}$). However, this can contribute to many interesting phenomena, as mentioned before. Nevertheless, the impact might be small. Moreover, we have considered the temperature evolution in calculating thermoelectric coefficients and estimating induced fields. Evolution contributes significantly at the early stage, enhancing the induced electric field compared to the static picture. In peripheral collisions, we consider an external time-varying magnetic field. The induced electric field is further enhanced by the external magnetic field. In other words, the magneto-thermoelectric effect induces a stronger electric field than the thermoelectric effect. The impact of the magnitude of the initial magnetic field and its decay parameter are also studied. At the early time of evolution, the magnetic decay parameter has a substantial impact; however, the strength of the initial magnetic field has a modest effect. Furthermore, we have accounted quantum modification of particles’ energy state due to Landau quantization in the presence of the magnetic field. Considering RHICs, the external magnetic field is taken along the $y$ direction for the numerical estimation, for which the induced field becomes anisotropic in the transverse plane. However, we have also generalized the calculations, where the direction of the magnetic field was not preferred, and we obtained the thermoelectric coefficient matrix for the same.
@@
Acknowledgements.
K.S. acknowledges the doctoral fellowship from UGC, Government of India. J.D. and R.S. gratefully acknowledge the DAE-DST, Govt. of India funding under the mega-science project – “Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123).
@@ A.1 Case-I (B = 0)
To find $\delta f_{i}$ in Eq. (15), we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA). In the presence of temperature gradients, the BTE under RTA can be expressed as [15]
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. $\vec{E}$ is an external electric field. We can solve Eq. (A.1) using Eqs. (18) and (19), and get the expression of $\alpha_{j}$’s. Here, the first term on left-hand side (lhs) of the Eq. (A.1) becomes,
$-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}{\dot{\vec{E}}}+\dot{\alpha_{2}}{\vec{\nabla}T}$
$+\alpha_{2}{\vec{\nabla}\dot{T}}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.2)
The second term in the lhs of the Boltzmann equation leads to,
$\frac{\partial f_{i}}{\partial x_{i}}$ $=\frac{\partial}{\partial x_{i}}(f^{0}_{i}+\delta f_{i})=\frac{\partial}{\partial x_{i}}(f^{0}_{i})+\frac{\partial f_{i}}{\partial x_{i}}(\delta f_{i})$
$=\frac{\partial}{\partial x_{i}}\Big{(}\frac{1}{1\pm\exp{\big{(}\beta(\omega_{i}-{\rm b}_{i}\mu_{i})\big{)}}}\Big{)}+0$
@@
The third term in the lhs leads to,
@@
$-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\omega_{i}{\vec{v}_{i}}.\Big{(}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}{\dot{\vec{E}}}+\dot{\alpha_{2}}{\vec{\nabla}T}+\alpha_{2}{\vec{\nabla}\dot{T}}\Big{)}$
$\frac{\partial f^{0}_{i}}{\partial\omega_{i}}-(\omega_{i}-{\rm b}_{i}h)\frac{\vec{\nabla}T}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}$
$=-\frac{\omega_{i}}{\tau_{R}^{i}}\Big{(}\alpha_{1}\vec{E}+\alpha_{2}\vec{\nabla}T+\alpha_{3}\vec{\nabla}\dot{T}\Big{)}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.6)
@@
$\dot{\alpha_{1}}$ $=-\Bigg{(}\frac{1}{\tau_{R}^{i}}\alpha_{1}+\frac{q_{i}}{\omega_{i}}\Bigg{)},~{}~{}~{}\dot{\alpha_{2}}=-\bigg{(}\frac{1}{\tau_{R}^{i}}\alpha_{2}-\frac{\omega_{i}-{\rm b}_{i}h}{\omega_{i}T}\bigg{)},$
${\alpha_{3}}$ $=-{\tau_{R}^{i}}\alpha_{2},$ (A.7)
@@
From Eq. (A.12) and Eq. (A.8), we can see that $Y_{p}$ is a column matrix with the same coefficients as that of matrix X. Further, the differentiation of Eq. (A.12) with respect to time gives us $Y_{p}^{{}^{\prime}}$ = $Y^{{}^{\prime}}U+YU^{{}^{\prime}}$, where $Y^{{}^{\prime}}=AY$. Hence, $Y_{p}^{{}^{\prime}}$ = $AY_{p}+YU^{{}^{\prime}}$. Comparison of above equation with Eq. (A.8) tells us
@@
$\alpha_{1}=-\frac{q_{i}\tau_{R}^{i}}{\omega_{i}},~{}~{}~{}\alpha_{2}=\frac{\tau_{R}^{i}(\omega_{i}-{\rm b}_{i}h)}{T\omega_{i}},~{}~{}~{}\alpha_{3}=-\tau_{R}^{i}\alpha_{2}.$
@@
$\delta f_{i}$ $=-q_{i}\tau_{R}^{i}(\vec{v_{i}}.\vec{E})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\frac{(\omega_{i}-{\rm b}_{i}h)\tau_{R}^{i}}{T}\Big{[}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{T}\Big{)}$
$-\tau_{R}^{i}\Big{(}\vec{v_{i}}\cdot\vec{\nabla}{\dot{T}}\Big{)}\Big{]}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.18)
Here, $\pm$ indicates a positive sign for positively charged particles(antiparticles) and a negative sign for negatively charged particles (antiparticles).
@@ A.2 Case-II (B $\not$= 0)
To find the expression of $\delta f_{i}$ in Eq. (15), we solve the Boltzmann transport equation (BTE) with the help of relaxation time approximation (RTA). In the presence of an external electromagnetic field, BTE under RTA can be expressed as [15]
@@
where $\tau^{i}_{R}$ is the relaxation time of the particle. We can solve Eq. (A.19) using Eqs. (38) and (IV.1.1), and get the expression of $\alpha_{j}$’s. Here, the first term on the lhs of the Eq. (A.19) becomes,
$\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}\dot{{\vec{E}}}+\dot{\alpha_{3}}{\vec{\nabla}T}+{\alpha_{3}}{\vec{\nabla}\dot{T}}+\alpha_{5}({\vec{\nabla}T}\times\dot{{\vec{B}}})$
$+\dot{\alpha_{5}}({\vec{\nabla}T}\times{\vec{B}})+{\alpha_{5}}({\vec{\nabla}\dot{T}}\times{\vec{B}})+\alpha_{5}(\vec{\nabla}\dot{T}\times{\vec{B}})+\alpha_{8}({\vec{E}}\times\dot{{\vec{B}}})$
$+\dot{\alpha_{8}}({\vec{E}}\times{\vec{B}})+\alpha_{8}(\dot{\vec{E}}\times{\vec{B}})-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.20)
The second term in the lhs of the Boltzmann equation leads to,
$\frac{\partial f_{i}}{\partial x_{i}}$ $=\frac{\partial}{\partial x_{i}}(f^{0}_{i}+\delta f_{i})=\frac{\partial}{\partial x_{i}}(f^{0}_{i})+\frac{\partial f_{i}}{\partial x_{i}}(\delta f_{i})$
$=\frac{\partial}{\partial x_{i}}\Big{\{}\frac{1}{1\pm\exp{\big{(}\beta(\omega_{i}-{\rm b}_{i}\mu_{i})\big{)}}}\Big{\}}+0$
@@
The third term in the lhs leads to,
@@
$q_{i}\left(\vec{E}+\frac{\vec{k}_{i}}{\omega_{i}}\times\vec{B}\right)\cdot\frac{\partial f_{i}}{\partial\vec{k_{i}}}=q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\big{(}\vec{v}\times\vec{B}\big{)}$
$\big{\{}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})$
$+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})$
$+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}})\big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.23)
@@
$\omega_{i}{\vec{v}_{i}}.\Big{\{}\dot{\alpha_{1}}{\vec{E}}+\alpha_{1}\dot{{\vec{E}}}+\dot{\alpha_{3}}{\vec{\nabla}T}+{\alpha_{3}}{\vec{\nabla}\dot{T}}+\alpha_{5}({\vec{\nabla}T}\times\dot{{\vec{B}}})+\dot{\alpha_{5}}({\vec{\nabla}T}\times{\vec{B}})+{\alpha_{5}}({\vec{\nabla}\dot{T}}\times{\vec{B}})+\alpha_{5}(\vec{\nabla}\dot{T}\times{\vec{B}})+\alpha_{8}({\vec{E}}\times\dot{{\vec{B}}})$
$+\dot{\alpha_{8}}({\vec{E}}\times{\vec{B}})+\alpha_{8}(\dot{\vec{E}}\times{\vec{B}})-(\omega_{i}-{\rm b}_{i}\mu)\frac{\dot{T}}{T}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.-(\omega_{i}-{\rm b}_{i}h)\vec{v_{i}}\cdot\frac{\vec{\nabla}T}{T}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+q_{i}\vec{v_{i}}\cdot\vec{E}\frac{\partial f^{0}_{i}}{\partial\omega_{i}}+\big{\{}-\alpha_{1}q_{i}\vec{v_{i}}\cdot(\vec{E}\times\vec{B})$
$-\alpha_{2}q_{i}\vec{v_{i}}\cdot(\dot{\vec{E}}\times\vec{B})-\alpha_{3}q_{i}\vec{v_{i}}\cdot(\vec{\nabla}T\times\vec{B})-\alpha_{4}q_{i}\vec{v_{i}}\cdot(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{5}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\vec{\nabla T}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{\nabla}T)({\vec{B}}\cdot\vec{B})\}$
@@
$+\alpha_{8}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\vec{E}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{E})({\vec{B}}\cdot\vec{B})\}+\alpha_{9}q_{i}\{-(\vec{v_{i}}\cdot\dot{\vec{B}})(\vec{E}\cdot\vec{B})+(\vec{v_{i}}\cdot\vec{\dot{E}})({\vec{B}}\cdot\dot{\vec{B}})\}+\alpha_{9}q_{i}\{-(\vec{v_{i}}\cdot\dot{\vec{B}})(\vec{E}\cdot\vec{B})$
$+(\vec{v_{i}}\cdot\dot{\vec{E}})({\vec{B}}\cdot\dot{\vec{B}})\}+\alpha_{10}q_{i}\{-(\vec{v_{i}}\cdot\vec{B})(\dot{\vec{E}}\cdot\vec{B})+(\vec{v_{i}}\cdot\dot{\vec{E}})({\vec{B}}\cdot\vec{B})\}\Big{\}}\frac{\partial f^{0}_{i}}{\partial\omega_{i}},$ (A.24)
@@
$=-\frac{\omega_{i}}{\tau_{R}^{i}}\alpha_{1}\vec{E}+\alpha_{2}\dot{\vec{E}}+\alpha_{3}\vec{\nabla}T+\alpha_{4}\vec{\nabla}\dot{T}+\alpha_{5}(\vec{\nabla}T\times\vec{B})+\alpha_{6}(\vec{\nabla}T\times\dot{\vec{B}})+\alpha_{7}(\vec{\nabla}\dot{T}\times\vec{B})+\alpha_{8}(\vec{E}\times{\vec{B}})+\alpha_{9}(\vec{E}\times{\dot{\vec{B}}})$
$+\alpha_{10}(\dot{\vec{E}}\times{\vec{B}})\frac{\partial f^{0}_{i}}{\partial\omega_{i}}.$ (A.25)
@@
Here, $\dot{\alpha_{1}}$, $\dot{\alpha_{3}}$, $\dot{\alpha_{5}}$ and $\dot{\alpha_{8}}$ from Eq. (A.2) can be expressed in terms of matrix equation as
@@
$X$ $=\begin{pmatrix}\alpha_{1}\\ \alpha_{3}\\ \alpha_{5}\\ \alpha_{8}\end{pmatrix},~{}~{}~{}A=\begin{pmatrix}-\frac{1}{\tau_{R}^{i}}&0&0&-\frac{q_{i}F^{2}}{\omega_{i}}\\ 0&-\frac{1}{\tau_{R}^{i}}&-\frac{q_{i}F^{2}}{\omega_{i}}&0\\ 0&\frac{q_{i}}{\omega_{i}}&-\frac{1}{\tau_{R}^{i}}&0\\ \frac{q_{i}}{\omega_{i}}&0&0&-\frac{1}{\tau_{R}^{i}},\end{pmatrix},$
$G$ $=\begin{pmatrix}-\frac{q_{i}}{\omega_{i}}\\ \frac{\omega_{i}-{\rm b}_{i}h}{T\omega_{i}}\\ 0\\ 0\end{pmatrix},$